OpenStudy (itsmichelle29):

i give medals g(x) = x2 +6x + 1 solve

OpenStudy (itsmichelle29):

in quadratic formula

OpenStudy (itsmichelle29):

thanks

OpenStudy (anonymous):

\[x=-3\pm 2 \sqrt{2} \]

OpenStudy (itsmichelle29):

and what are the steps

OpenStudy (freckles):

what are you solving for?

OpenStudy (itsmichelle29):

help plz

OpenStudy (itsmichelle29):

g(x) = x2 +6x + 1

OpenStudy (freckles):

but what do we solve?

OpenStudy (freckles):

are you trying to solve g(x)=0 for x?

OpenStudy (freckles):

are you trying ting solve g(x)=1 ?

OpenStudy (itsmichelle29):

in quadratic formula

OpenStudy (freckles):

so on...

OpenStudy (itsmichelle29):

i know robertobey is the right answer i just need the steps

OpenStudy (itsmichelle29):

can you help me on that

OpenStudy (freckles):

so are you solving x^2+6x+1=0 ?

OpenStudy (itsmichelle29):

yup

OpenStudy (freckles):

ok I didn't see this equation anywhere when you said solve so you can understand my confusion on what we want to solve

OpenStudy (itsmichelle29):

yes

OpenStudy (freckles):

anyways first compare x^2+6x+1=0 to ax^2+bx+c=0 identify a,b, and c?

OpenStudy (itsmichelle29):

okay so i dont know what to do after x= -6 plus and minus root 32 all over 2

OpenStudy (itsmichelle29):